OpenStudy (anonymous):

A trip to portland, oregon from Boston will take 7 3/4 hours. Assuming we are 2/3 of the way there, how much longer in hours will the trip take?

OpenStudy (stormfire1):

\[7 \frac{3}{4}-(\frac{2}{3}*7\frac{3}{4})=2~hrs~35~minutes\]